OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? (2, 1) (1, 1) (0, 1)

OpenStudy (anonymous):

(1, 1)

OpenStudy (anonymous):

thank you ademm(:

OpenStudy (anonymous):

Well let's first get the equation. We use point-slope. y-y1=m(x-x1). y+1=2(x-0) = y+1=2x+0 Now we subtract both sides by 1 y=2x-1. Now all we have to do is put in those 3 values and see which one equals out. Once you do, the only that works is (1,1) and that's your answer.

OpenStudy (anonymous):

thankk you loujoelou